# Computing almost commuting bases of ODOs and Gelfand-Dikey hierarchies

### v1

@@ 4.2 Solving by integration in closed form
Let $L$ be the formal differential operator of order $n\geq 2$ in $\mathbb{Q}\{U\}[\partial]$ and in normal form as in (7). At this point, for a fixed $m\geq 2$, we would like to solve the triangular differential system $\mathbb{E}_{m,n}$ with respect to the set of variables $Y_{m}=\{y_{2},\dots,y_{m}\}$. Recall that $Y_{i}=\{y_{2},\dots,y_{i}\}$, $12\leq j\leq m+$ 1 and define $Y_{1}:=\emptyset$. By Theorem 4.5 the equations of $\mathbb{E}_{m,n}$ have the form
@@
with $e_{m,i-2}\in\mathbb{Q}\{U\}\{Y_{i-2}\}$ homogeneous of weight $i$. Furthermore, at each step, the new variable that we solve for, $y_{i-1}$, appears differentiated only once.
@@
Observe that $\partial^{-1}e_{m,i}(q_{2},\ldots,q_{i-1})$ is well defined by Lemma (4.6).
@@ 4       $Z\leftarrow Z\cup\{y_{i}=q_{i}\}$;
In step 4 of Algorithm 2 the instruction $\texttt{INTEG}(e_{m,i})$ indicates the computation of $\partial^{-1}e_{m,i}(q_{2},\ldots,q_{i-1})$. For this purpose we use the canonical decomposition of differential polynomials in one differential variable introduced in  [1]. It was extended in the context of differential algebra in the article [2], for partial derivations and general rankings.
Namely, given a differential polynomial $F(z)\in\mathbb{Q}\{z\}$, an algorithm was provided in  [1] to write this polynomial as $F(z)=\partial(A(z))+B(z)$, where $A(z)\in\mathbb{Q}\{z\}$ and $B(z)$ is a functional polynomial as defined in [1, 2], whose monomials are not integrable, they contain only non-linear monomials in their highest derivative. This algorithm works for multivariate differential polynomials in $\mathbb{Q}\{U\}$ iterating the algorithm over each of the differential variables $u_{2}>u_{3}>\cdots>u_{n}$ and collecting the final result.
These methods in [1, 2] provide a criteria for integrability (i.e., $F(z)$ is integrable if and only if $B(z)=0$). We have implemented this idea to compute antiderivatives of differential polynomials in our SageMath package, dalgebra (see Section 6).
@@ 5 Computing integrable hierarchies
where $A_{m}$ is an almost commuting differential operator of order $m$. Observe that $[A_{m},L]$ is a multiplication operator.
@@
Let us consider a differential field of constants $(\mathbf{C},\partial)$ and assume that $\mathbf{C}$ has zero characteristic, thus $\mathbf{C}$ contains $\mathbb{Q}$. The Lax equations in this case are
@@
More precisely, let us consider a fixed $m\in\mathbb{N}$ and the operator $P_{m}$ in the basis obtained with our method and $Z=\{q_{2},\ldots,q_{n}\}$ be the elements that allow $\epsilon_{Z}(\tilde{P}_{m})=P_{m}$. Since $P_{m}$ almost commutes with $L$, then $[L,P_{m}]$ is a differential operator with order $n-2$. As described in Theorem 4.1, we obtain:
$[L,P_{m}]=[L,\epsilon_{Z}(\tilde{P}_{m})]=\sum_{i=2}^{n}E_{m,n-i}(U,Z)\partial^{n-i}.$
Let us denote the specialization of $-E_{m,n-i}(U,Y)$ by $\epsilon_{Z}$ by
@@
Consequently, the Lax equations (25) provide a system of partial differential equations defined by non linear differential polynomials in the set of differential variables $U$, obtained from the coefficient of $\partial^{n-i}$ in $[A_{m},L]$
@@
For a fixed $n$, the family (27) is called the Gelfand-Dickey hierarchy of $L_{n}$, see [5, 6, 21] and the references therein. Some of these hierarchies have been given specific names, for instance, for $n=2$, the Gelfand-Dikey hierarchy of $L_{2}=\partial^{2}+u_{2}$ is the Korteweg-de Vries (KdV) hierarchy
@@
For $n=3$, the Gelfand-Dikey hierarchy of $L_{3}=\partial^{3}+u_{2}\partial+u_{3}$ is
$u_{i,t}=H_{m,3-i}(U)+\sum_{j=1}^{m-1}c_{m,j}H_{j,3-i}(U)\ ,\textrm{for }\ i=2,\dots,3,\,\,\,m\geq 2.$ (29)
As we will see in Section 5.1, these systems can be identified with the Bousinesq systems [3].
We obtain from (27) the stationary Gelfand-Dickey hierarchy of $L_{n}$ considering $t=0$. These nonlinear differential polynomials can be seen as conditions over the differential variables $U$ for an almost commuting operator $A_{m}$ to commute with $L_{n}$. See [6, 21] and the references therein.
5.1 Third order operators and their hierarchy
@@
For each $m\geq 1$, it holds
$[L_{3},P_{m}]=H_{m,1}(U)\partial+H_{m,0}(U),$
where $H_{m,0}(U)$ and $H_{m,1}(U)$ are differential polynomials in $\mathcal{R}$ that determine all the equations of the Gelfand-Dikey hierarchy for $L_{3}$. We will compute next the equations (27) for $2\leq m\leq 5$.
Observe that $P_{1}=\partial$, thus $[L_{3},P_{1}]=-u_{2}^{\prime}\partial-u_{3}^{\prime}$. Hence we have
@@
$[L_{3},\tilde{P}_{2}]=E_{2,2}\partial^{2}+E_{2,1}\partial+E_{2,0},$
@@
$\begin{array}[]{rcl}E_{2,2}&=&\mathbf{3y_{2}^{\prime}-2u_{2}^{\prime}},\\ E_{2,1}&=&3y_{2}^{\prime\prime}-u_{2}^{\prime\prime}-2u_{3}^{\prime},\\ E_{2,0}&=&y^{\prime\prime\prime}+u_{2}y_{2}-u_{3}^{\prime\prime}.\\ \end{array}$
In order to obtain the almost commuting operator $P_{2}$, we force the coefficient of $\partial^{2}$ of to be zero. We compute $q_{2}$ such that $E_{2,2}(u_{2},q_{2})=0$ and obtain
@@
$[L_{3},\tilde{P}_{4}]=E_{4,4}\partial^{4}+E_{4,3}\partial^{3}+E_{4,2}\partial^{2}+E_{4,1}\partial+E_{4,0},$
@@
$\hss\begin{array}[]{rcl}E_{4,4}&=&\mathbf{3y_{2}^{\prime}-4u_{2}^{\prime}},\\ E_{4,3}&=&\mathbf{3y_{3}^{\prime}+3y_{2}^{\prime\prime}-6u_{2}^{\prime\prime}-4u_{3}^{\prime}},\\[2.5pt] E_{4,2}&=&\mathbf{3y_{4}^{\prime}+3y_{3}^{\prime\prime}+y_{2}^{\prime\prime\prime}+u_{2}y_{2}^{\prime}-2u_{2}^{\prime}y_{2}-4u_{2}^{\prime\prime\prime}-6u_{3}^{\prime\prime}},\\[2.5pt] E_{4,1}&=&3y_{4}^{\prime\prime}+y_{3}^{\prime\prime\prime}+u_{2}y_{3}^{\prime}-u_{2}^{\prime\prime}y_{2}-u_{2}^{\prime}y_{3}-2u_{3}^{\prime}y_{2}-u_{2}^{(4)}-4u_{3}^{\prime\prime\prime},\\[2.5pt] E_{4,0}&=&y_{4}^{\prime\prime\prime}+u_{2}y_{4}^{\prime}-u_{3}^{\prime\prime}y_{2}-u_{3}^{\prime}y_{3}-u_{3}^{(4)}.\end{array}\hss$
@@
$P_{4}=\mathcal{E}_{Z}(\tilde{P}_{4})=\partial^{4}+\frac{4}{3}u_{2}\partial^{2}+\left(\frac{2}{3}u_{2}^{\prime}+\frac{4}{3}u_{3}\right)\partial+\left(\frac{2}{9}u_{2}^{\prime\prime}+\frac{2}{3}u_{3}^{\prime}+\frac{2}{9}u_{2}^{2}\right).$
@@
$-H_{4,0}(u_{2},u_{3})=$ $\frac{4}{9}u_{2}^{\prime}u_{2}^{2}+\frac{2}{3}u_{2}^{\prime\prime\prime}u_{2}+\frac{4}{3}u_{2}^{\prime\prime}u_{2}^{\prime}-\frac{2}{3}u_{2}^{\prime}u_{3}^{\prime}-\frac{2}{3}u_{2}u_{3}^{\prime\prime}-\frac{4}{3}u_{3}^{\prime}u_{3}+\frac{2}{9}u_{2}^{(5)}$
$-\frac{1}{3}u_{3}^{(4)}$
@@
$[L_{3},\tilde{P}_{5}]=E_{5,5}\partial^{5}+E_{5,4}\partial^{4}+E_{5,3}\partial^{3}+E_{5,2}\partial^{2}+E_{5,1}\partial+E_{5,0},$
@@
$\begin{array}[]{ll}E_{5,5}=&{\bf 3y_{2}^{\prime}-5u_{2}^{\prime}},\\ E_{5,4}=&{\bf 3y_{3}^{\prime}+3y_{2}^{\prime\prime}-10u_{2}^{\prime\prime}-5u_{3}^{\prime}},\\ E_{5,3}=&{\bf 3y_{4}^{\prime}+3y_{3}^{\prime\prime}+y_{2}^{\prime\prime\prime}+u_{2}y_{2}^{\prime}-3u_{2}^{\prime}y_{2}-10u_{2}^{\prime\prime\prime}-10u_{3}^{\prime\prime}},\\ E_{5,2}=&{\bf 3y_{5}^{\prime}+3y_{4}^{\prime\prime}+y_{3}^{\prime\prime\prime}+u_{2}y_{3}^{\prime}-3u_{2}^{\prime\prime}y_{2}-2u_{2}^{\prime}y_{3}-3u_{3}^{\prime}y_{2}-5u_{2}^{(4)}-10u_{3}^{\prime\prime\prime}},\\ E_{5,1}=&3y_{5}^{\prime\prime}+y_{4}^{\prime\prime\prime}+u_{2}y_{4}^{\prime}-u_{2}^{\prime\prime\prime}y_{2}-u_{2}^{\prime\prime}y_{3}-u_{2}^{\prime}y_{4}-3u_{3}^{\prime\prime}y_{2}-2u_{3}^{\prime}y_{3}-u_{2}^{(5)}-5u_{3}^{(4)},\\ E_{5,0}=&u_{2}y_{5}^{\prime}-u_{3}^{\prime\prime\prime}y_{2}-u_{3}^{\prime\prime}y_{3}-u_{3}^{\prime}y_{4}-u_{3}^{(5)}+y_{5}^{\prime\prime\prime}.\end{array}$
@@
$P_{5}=\mathcal{E}_{Z}(\tilde{P}_{5})=$ $\partial^{5}+\frac{5}{3}u_{2}\partial^{3}+\left(\frac{5}{3}u_{2}^{\prime}+\frac{5}{3}u_{3}\right)\partial^{2}+\left(\frac{10}{9}u_{2}^{\prime\prime}+\frac{5}{3}u_{3}^{\prime}+\frac{5}{9}u_{2}^{2}\right)\partial$
$+\left(\frac{10}{9}u_{3}^{\prime\prime}+\frac{10}{9}u_{2}u_{3}\right).$
@@
$\left\{\begin{array}[]{rcl}-H_{5,0}(u_{2},u_{3})&=&\displaystyle\frac{10}{9}u_{2}^{\prime}u_{2}u_{3}+\frac{5}{9}u_{2}^{2}u_{3}^{\prime}+\frac{10}{9}u_{2}^{\prime\prime\prime}u_{3}+\frac{20}{9}u_{2}^{\prime\prime}u_{3}^{\prime}\\[10.00002pt] &&\displaystyle+\frac{5}{3}u_{2}^{\prime}u_{3}^{\prime\prime}+\frac{5}{9}u_{2}u_{3}^{\prime\prime\prime}-\frac{5}{3}u_{3}^{\prime\prime}u_{3}-\frac{5}{3}u_{3}^{\prime 2}+\frac{1}{9}u_{3}^{(5)}\\[15.00002pt] -H_{5,1}(u_{2},u_{3})&=&\displaystyle\frac{5}{9}u_{2}^{\prime}u_{2}^{2}+\frac{5}{9}u_{2}^{\prime\prime\prime}u_{2}+\frac{5}{9}u_{2}^{\prime\prime}u_{2}^{\prime}+\frac{5}{3}u_{2}^{\prime\prime}u_{3}+\frac{5}{3}u_{2}^{\prime}u_{3}^{\prime}\\[10.00002pt] &&\displaystyle-\frac{10}{3}u_{3}^{\prime}u_{3}+\frac{1}{9}u_{2}^{(5)}\end{array}\right.$
5.2 Fith order operators and their hierarchy
Let us consider the third order operator in $\mathcal{R}[\partial]$ with $\mathcal{R}=\mathbf{C}\{U\}$, with $U=\{u_{2},u_{2},u_{4},u_{5}\}$ given by

### v2

@@ 4.2 Solving by integration in closed form
Let $L$ be the formal differential operator of order $n\geq 2$ in $\mathbb{Q}\{U\}[\partial]$ and in normal form as in (7). At this point, for a fixed $m\geq 2$, we would like to solve the triangular differential system $\mathbb{E}_{m,n}$ with respect to the set of variables $Y_{m}=\{y_{2},\dots,y_{m}\}$. Recall that $Y_{i}=\{y_{2},\dots,y_{i}\}$, $2\leq j\leq m+1$ and define $Y_{1}:=\emptyset$. By Theorem 4.6 the equations of $\mathbb{E}_{m,n}$ have the form
@@
with $e_{m,i-2}\in\mathbb{Q}\{U\}\{Y_{i-2}\}$ homogeneous of weight $i$. Furthermore, at each step, the new variable $y_{i-1}$, that we solve for, appears differentiated only once.
@@
Observe that $\partial^{-1}e_{m,i}(q_{2},\ldots,q_{i-1})$ is well defined by Lemma 4.7.
@@ 4  $Z\leftarrow Z\cup\{y_{i}=q_{i}\}$;
In step 4 of Algorithm 2 the instruction $\texttt{INTEG}(e_{m,i})$ indicates the computation of $\partial^{-1}e_{m,i}(q_{2},\ldots,q_{i-1})$. For this purpose we use the canonical decomposition of differential polynomials in one differential variable introduced in  [14]. This method was extended in the context of differential algebra in the article [15], for partial derivations and general rankings.
Namely, given a differential polynomial $F(z)\in\mathbb{Q}\{z\}$, an algorithm was provided in  [14] to write this polynomial in a unique form $F(z)=\partial(A(z))+B(z)$, where $B(z)$ is not in the image of $\partial$ [14, 15], they contain only non-linear monomials in their highest derivative. This algorithm works for multivariate differential polynomials in $\mathbb{Q}\{U\}$ by iterating the algorithm over each of the differential variables $u_{2}>u_{3}>\cdots>u_{n}$ and collecting the final result.
The methods in [14, 15] provide a criteria to detect total derivatives (i.e., $F(z)$ is a total derivative if and only if $B(z)=0$). We have implemented this method to compute antiderivatives of differential polynomials in our SageMath package, dalgebra (see Section 6).
@@ 5 Computing integrable hierarchies
where $A_{m}$ is an almost commuting differential operator of order $m$. Observe that $A_{m},L]$ is a multiplication operator.
@@
Let us consider a field of constants $(\mathbf{C},\partial)$ and assume that $\mathbf{C}$ has zero characteristic, thus $\mathbf{C}$ contains $\mathbb{Q}$. The Lax equations in this case are
@@
More precisely, let us consider a fixed $m\in\mathbb{N}$ and the operator $P_{m}$ in the almost commuting basis obtained with Algorithm  1 and $Z=\{q_{2},\ldots,q_{n}\}$ be the elements that allow $\epsilon_{Z}(\tilde{P}_{m})=P_{m}$. Since $P_{m}$ almost commutes with $L$, then $L,P_{m}]$ is a differential operator with order n2n-2. As described in Theorem 4.2, we obtain:
$L,P_{m}]=[L,\epsilon_{Z}(\tilde{P}_{m})]=\sum_{i=2}^{n}E_{m,n-i}(U,Z)\partial^{n-i}.$
Let us denote the specialization of $E_{m,n-i}(U,Y)$ by $\epsilon_{Z}$ as
@@
Consequently, the Lax equations (26) provide a system of partial differential equations defined by non linear differential polynomials in the set of differential variables $U$, obtained from the coefficient of $\partial^{n-i}$ in $A_{m},L]$
@@
For a fixed $n$, the family (28) is called the Gelfand-Dickey (GD) hierarchy of $L_{n}$, see [2, 3, 1] and references therein. Some of these hierarchies have been given specific names, for instance, for $n=2$, the Gelfand-Dickey hierarchy of $L_{2}=\partial^{2}+u_{2}$ is the Korteweg-de Vries (KdV) hierarchy
@@
For $n=3$, the Boussinesq hierarchy is obtained, see Section 5.1.
Other methods have been developed to compute the GD hierarchies, although normally they are particular methods for a fixed level, a fixed value of $n$. It is well known that the KdV hierarchy can be computed using recursion operators. We can rewrite (29) as
$u_{2,t}=\text{kdv}_{m}(u_{2})+\sum_{j=1}^{m-1}c_{m,j}\text{kdv}_{j}(u_{2})\ ,\,\,\,m\geq 2.$ (30)
where $\{\text{kdv}_{j}(u_{2})\}_{j\geq 0}$ is computed using the recursion operator, see [28] based on [7],
$\mathcal{R}=-\frac{1}{4}\partial^{2}+u_{2}+\frac{1}{2}u_{2}^{\prime}\partial^{-1},$
defining the recursion
$\text{kdv}_{0}:=u^{\prime},\,\,\,\text{kdv}_{n}:=\mathcal{R}(\text{kdv}_{n-1}),\mbox{ for }n\geq 1.$ (31)
We obtain from (28) the stationary Gelfand-Dickey hierarchy of $L_{n}$ considering the differential variables in $U$ as constants with respect to $t$. These nonlinear differential polynomials can be seen as conditions over the differential variables $U$ for an almost commuting operator $A_{m}$ to commute with $L_{n}$. See [3, 1] and the references therein.
5.1 Third order operators and their GD hierarchy
@@
The Gelfand-Dickey hierarchy of $L_{3}$ consists of systems of two nonlinear equations. For each $m\geq 2$
$u_{2,t}=H_{m,1}(u_{2},u_{3})+\sum_{j=1}^{m-1}c_{m,j}H_{j,1}(u_{2},u_{3}),$ (32)
$u_{3,t}=H_{m,0}(u_{2},u_{3})+\sum_{j=1}^{m-1}c_{m,j}H_{j,0}(u_{2},u_{3}).$ (33)
@@
For each $m\geq 1$, it holds that
$P_{m},L_{3}]=H_{m,1}(U)\partial+H_{m,0}(U),$
where $H_{m,0}(U)$ and $H_{m,1}(U)$ are differential polynomials in $\mathcal{R}$ that determine all the equations of the Gelfand-Dickey hierarchy for $L_{3}$. We will compute next the equations (32) for $2\leq m\leq 5$. Recall that $P_{1}=\partial$ by Example (3.9), thus $L_{3},P_{1}]=-u_{2}^{\prime}\partial-u_{3}^{\prime}$. Hence we have
@@
$L_{3},\tilde{P}_{2}]=E_{2,2}\partial^{2}+E_{2,1}\partial+E_{2,0},$
@@
$E_{2,2}$ $=\mathbf{3y_{2}^{\prime}-2u_{2}^{\prime}},$
$E_{2,1}$ $=3y_{2}^{\prime\prime}-u_{2}^{\prime\prime}-2u_{3}^{\prime},$
$E_{2,0}$ $=y_{2}^{\prime\prime\prime}+u_{2}y_{2}-u_{3}^{\prime\prime}.$
In order to obtain the almost commuting operator $P_{2}$, we force the coefficient of $\partial^{2}$ to be zero. We compute $q_{2}$ such that $E_{2,2}(u_{2},q_{2})=0$ and obtain
@@
Recursive formulas were given in  [9] to provide an elementary algebraic approach for the entire hierarchy of Boussinesq systems. The recursion operator for (35) can be found using the technique given in [13] or through liftings of known recursion operators for different forms of the Boussinesq equation. It is not the goal of this paper to explain how all these forms of the Boussinesq systems can be identified.
@@
$L_{3},\tilde{P}_{4}]=E_{4,4}\partial^{4}+E_{4,3}\partial^{3}+E_{4,2}\partial^{2}+E_{4,1}\partial+E_{4,0},$
@@
$E_{4,4}$ $=\mathbf{3y_{2}^{\prime}-4u_{2}^{\prime}},$
$E_{4,3}$ $=\mathbf{3y_{3}^{\prime}+3y_{2}^{\prime\prime}-6u_{2}^{\prime\prime}-4u_{3}^{\prime}},$
$E_{4,2}$ $=\mathbf{3y_{4}^{\prime}+3y_{3}^{\prime\prime}+y_{2}^{\prime\prime\prime}+u_{2}y_{2}^{\prime}-2u_{2}^{\prime}y_{2}-4u_{2}^{\prime\prime\prime}-6u_{3}^{\prime\prime}},$
$E_{4,1}$ $=3y_{4}^{\prime\prime}+y_{3}^{\prime\prime\prime}+u_{2}y_{3}^{\prime}-u_{2}^{\prime\prime}y_{2}-u_{2}^{\prime}y_{3}-2u_{3}^{\prime}y_{2}-u_{2}^{(4)}-4u_{3}^{\prime\prime\prime},$
$E_{4,0}$ $=y_{4}^{\prime\prime\prime}+u_{2}y_{4}^{\prime}-u_{3}^{\prime\prime}y_{2}-u_{3}^{\prime}y_{3}-u_{3}^{(4)}.$
@@
$P_{4}=\epsilon_{Z}(\tilde{P}_{4})=\partial^{4}+\frac{4}{3}u_{2}\partial^{2}+\left(\frac{2}{3}u_{2}^{\prime}+\frac{4}{3}u_{3}\right)\partial+\left(\frac{2}{9}u_{2}^{\prime\prime}+\frac{2}{3}u_{3}^{\prime}+\frac{2}{9}u_{2}^{2}\right).$
@@
$-H_{4,0}(u_{2},u_{3})$ $=\frac{4}{9}u_{2}^{\prime}u_{2}^{2}+\frac{2}{3}u_{2}^{\prime\prime\prime}u_{2}+\frac{4}{3}u_{2}^{\prime\prime}u_{2}^{\prime}-\frac{2}{3}u_{2}^{\prime}u_{3}^{\prime}-\frac{2}{3}u_{2}u_{3}^{\prime\prime}-\frac{4}{3}u_{3}^{\prime}u_{3}$
$\quad{}+\frac{2}{9}u_{2}^{(5)}-\frac{1}{3}u_{3}^{(4)},$
@@
$L_{3},\tilde{P}_{5}]=E_{5,5}\partial^{5}+E_{5,4}\partial^{4}+E_{5,3}\partial^{3}+E_{5,2}\partial^{2}+E_{5,1}\partial+E_{5,0},$
@@
$E_{5,5}$ $={\bf 3y_{2}^{\prime}-5u_{2}^{\prime}},$
$E_{5,4}$ $={\bf 3y_{3}^{\prime}+3y_{2}^{\prime\prime}-10u_{2}^{\prime\prime}-5u_{3}^{\prime}},$
$E_{5,3}$ $={\bf 3y_{4}^{\prime}+3y_{3}^{\prime\prime}+y_{2}^{\prime\prime\prime}+u_{2}y_{2}^{\prime}-3u_{2}^{\prime}y_{2}-10u_{2}^{\prime\prime\prime}-10u_{3}^{\prime\prime}},$
$E_{5,2}$ $={\bf 3y_{5}^{\prime}+3y_{4}^{\prime\prime}+y_{3}^{\prime\prime\prime}+u_{2}y_{3}^{\prime}-3u_{2}^{\prime\prime}y_{2}-2u_{2}^{\prime}y_{3}-3u_{3}^{\prime}y_{2}-5u_{2}^{(4)}-10u_{3}^{\prime\prime\prime}},$
$E_{5,1}$ $=3y_{5}^{\prime\prime}+y_{4}^{\prime\prime\prime}+u_{2}y_{4}^{\prime}-u_{2}^{\prime\prime\prime}y_{2}-u_{2}^{\prime\prime}y_{3}-u_{2}^{\prime}y_{4}-3u_{3}^{\prime\prime}y_{2}-2u_{3}^{\prime}y_{3}-u_{2}^{(5)}-5u_{3}^{(4)},$
$E_{5,0}$ $=u_{2}y_{5}^{\prime}-u_{3}^{\prime\prime\prime}y_{2}-u_{3}^{\prime\prime}y_{3}-u_{3}^{\prime}y_{4}-u_{3}^{(5)}+y_{5}^{\prime\prime\prime}.$
@@
$P_{5}=\epsilon_{Z}(\tilde{P}_{5})$ $=\partial^{5}+\frac{5}{3}u_{2}\partial^{3}+\left(\frac{5}{3}u_{2}^{\prime}+\frac{5}{3}u_{3}\right)\partial^{2}+\left(\frac{10}{9}u_{2}^{\prime\prime}+\frac{5}{3}u_{3}^{\prime}+\frac{5}{9}u_{2}^{2}\right)\partial$
$\quad{}+\left(\frac{10}{9}u_{3}^{\prime\prime}+\frac{10}{9}u_{2}u_{3}\right).$
@@
$\left\{\begin{aligned} -H_{5,0}(u_{2},u_{3})&=\frac{10}{9}u_{2}^{\prime}u_{2}u_{3}+\frac{5}{9}u_{2}^{2}u_{3}^{\prime}+\frac{10}{9}u_{2}^{\prime\prime\prime}u_{3}+\frac{20}{9}u_{2}^{\prime\prime}u_{3}^{\prime}\\[10.00002pt] &\quad{}+\frac{5}{3}u_{2}^{\prime}u_{3}^{\prime\prime}+\frac{5}{9}u_{2}u_{3}^{\prime\prime\prime}-\frac{5}{3}u_{3}^{\prime\prime}u_{3}-\frac{5}{3}u_{3}^{\prime 2}+\frac{1}{9}u_{3}^{(5)},\\[15.00002pt] -H_{5,1}(u_{2},u_{3})&=\frac{5}{9}u_{2}^{\prime}u_{2}^{2}+\frac{5}{9}u_{2}^{\prime\prime\prime}u_{2}+\frac{5}{9}u_{2}^{\prime\prime}u_{2}^{\prime}+\frac{5}{3}u_{2}^{\prime\prime}u_{3}+\frac{5}{3}u_{2}^{\prime}u_{3}^{\prime}\\[10.00002pt] &\quad{}-\frac{10}{3}u_{3}^{\prime}u_{3}+\frac{1}{9}u_{2}^{(5)}.\end{aligned}\right$.
More results on the computation of the almost commuting basis and the differential polynomials defining the GD hierarchy for $n=3$ and $m\geq 2$ can be found in the repository da_wilson, see Section 6.
5.2 Fifth order operators and their GD hierarchy
Let us consider the third order operator in $\mathcal{R}[\partial]$ with $\mathcal{R}=\mathbf{C}\{U\}$, with $U=\{u_{2},{u_{3}},u_{4},u_{5}\}$ given by